OpenStudy (anonymous):

Decide whether the ordered pair is a solution of the given system. (-1, 6)

OpenStudy (anonymous):

(-1, 6) x + y = 7 x - y = -5

OpenStudy (anonymous):

sorry left out some part of the question but i think it is

OpenStudy (anonymous):

plug it in, do you get -5?

OpenStudy (anonymous):

x=-1, and y=6, from the ordered pair. so just plug those into your given system of equations.

OpenStudy (anonymous):

(x,y)=(-1,6)

OpenStudy (jwheele1):

X , Y (-1, 6) x + y = 7 x - y = -5 (-1) + 6 = 7(-1) -6 = -5 5 = -7 -6 = -5 5 = -13 = -5 No of those numbers equal each other. statement is false.

OpenStudy (anonymous):

Yo jwheele1, don't solve it for the person. Help them. that is what this website is about, I think

OpenStudy (anonymous):

no solution

OpenStudy (jwheele1):

I usually solve it but show the work So that they can see each step.

OpenStudy (jwheele1):

On bigger more complex problems i do break it up and ask questions

OpenStudy (anonymous):

Aright my bad

OpenStudy (anonymous):

what I'm I right no solution?

OpenStudy (anonymous):

@Esmy The ordered pair is not a solution of the given system. The statement is false.

OpenStudy (anonymous):

okay thank you ... :)

OpenStudy (jwheele1):

The answer is not a number or no solution....the anser is either true r false....in this case false

OpenStudy (jwheele1):

or..."no, theordered pair does not provide a solution to the given system."